OpenStudy (anonymous):

The Area of a circle increases at a rate of 1 cm^2/s. How fast is the radius changing when the circumference is 2 cm?

OpenStudy (anonymous):

This is a related rates question. Do you know the first step to tackling these?

OpenStudy (anonymous):

Yes. But i dont know how to apply the formulas

OpenStudy (anonymous):

A=pi r^2 and C=2pi r

OpenStudy (anonymous):

Ok, and they give you the rate at which it is increasing this is the differential dA/dt=1

OpenStudy (anonymous):

i dont get how you can find all the variables

OpenStudy (anonymous):

And now we need to find the radius first off to plug it into our final formula. Alright we know that C=2 pi r and they tell us our C is 2 so it will become \[\Large 2=2\pi~r\] Solve for r first off

OpenStudy (anonymous):

r= 2/2pi

OpenStudy (anonymous):

Alright and that simplifies to 1/pi now we can move on to differentiate the Area (since that's what we're working with) and find it with respect to t (time) do you know how to do that? It's kind of like implicit differentiation with the circle equation \[\Huge x^2+y^2=r^2\] where r is the radius and would be a constant.

OpenStudy (anonymous):

differentiate the equation?

OpenStudy (anonymous):

where did this eqtn come from ?

OpenStudy (anonymous):

but i think its 2x dx/dt +2y dy/dt = 1/pi^2

OpenStudy (anonymous):

No, no sorry I was just referecning to this equation

OpenStudy (anonymous):

I just meant for the area formula can you implicitly differentiate which I see you can and so we'd have \[\Huge A=\pi~r^2\] and we need to differentiate with respect to (time) and sorr @Jennings15 my computer restarted

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

dA/dt = 2pi r dr/dt

OpenStudy (anonymous):

r=1/pi

OpenStudy (anonymous):

Ok and the question is asking for how fast the radius is chaning, the rate of change of the radius (dr/dt) We know dA/dt to be 1 since they gave it to us (1cm^2) and we know r to be 1/pi so we just plug in and solve for dr/dt

OpenStudy (anonymous):

1cm^2/s *

OpenStudy (anonymous):

omg wow man . thats freaking amazing. let me solve this nw

OpenStudy (anonymous):

1=2pi/pi dr/dt 1=2dr/dt 1/2=dr/dt

OpenStudy (anonymous):

Anytime, I remember these, let me guess you're in Calc AB or some type of calculus class?

OpenStudy (anonymous):

preciate it bro , and yeah im in calculus/analytic gemotery 1 .

OpenStudy (anonymous):

And that looks about right, make sure to put your answer in the correct units so it's a length so it would just be 1.2cm/s

OpenStudy (anonymous):

1/2 cm/s*